OpenStudy (anonymous):

Natalie made a mistake in the following problem. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 6(4) ÷ 2 − 1 + 5 Line 2 6(4) ÷ 1 + 5 Line 3 24 ÷ 1 + 5 Line 4 24 + 5 Line 5 29

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6 * 4 = 24 24

OpenStudy (anonymous):

24/2 = 12

OpenStudy (anonymous):

12 - 1 = 11

OpenStudy (anonymous):

11 + 5 = 16

OpenStudy (anonymous):

y eah., that's going to be line 2 with the msistake.. sorry for the ytypos

OpenStudy (anonymous):

line 2 is wrong

OpenStudy (anonymous):

the answer would be 16

OpenStudy (anonymous):

I hope this helps!

OpenStudy (anonymous):

Did you get all of this down? Line 2 wron g and answer 16